# How Does Knowledge Selection Help Retrieval Augmented Generation?

### v2

@@ 6.2 Recommendations to Future Practitioners
Only when the knowledge recall score is high can a knowledge selector, which increases knowledge precision, be potentially helpful***In a length-constrained scenario, if we consider $k$ randomly sampled knowledge sentences as a baseline, rather than the “full knowledge” setting, then the knowledge selector is more likely to be helpful due to weaker baseline performance.. Moreover, if the recall is too low, the downstream generation performance may not see any benefits from a knowledge selector with middling performance — it may even be harmed; only when the knowledge selector’s precision is very high will the overall performance improve.
@@
Due to the high cost of larger-scale simulation experiments using API-based LLMs, we use only a subset of the WoW and HotpotQA for experiments. As a result, our data subset may contain some minor noise during knowledge sampling and LLM generation, and the contours in Figure 4 are not smooth. However, such noise is not likely to affect our conclusions. For a similar reason, we only chose three LLMs for our experiments. As a result, we may have missed out on more subtle phenomena that can only be seen from results on a larger number of generators.
@@ A.1 Implementation Details
We do not perform any fine-tuning or hyperparameter tuning. We set the LLMs’ temperature to 0 and ensure that our zero-shot generation prompts (Appendix A.2) are shorter than their maximum input lengths. We use the API services of OpenAI gpt-4o-mini-2024-07-18***https://platform.openai.com/docs/models/gpt-4o-mini, Together AI***https://docs.together.ai/docs/chat-models meta-llama/Meta-Llama-3.1-8B-Instruct-Turbo, and mistralai/Mistral-7B-Instruct-v0.1. We report results on the first 500 examples in the HotpotQA training set and the first 100 conversations (452 wizard utterances) from the “test seen” set of WoW. Our set of experiments cost about 50 USD from OpenAI and about 50 USD from Together.ai.
@@ A.6 Length-Constrained Knowledge Selection and Answer Generation
To study the effect of length-constrained knowledge inputs on answer generation performance, we limit the number of knowledge sentences to $k$ by randomly sub-sampling $k$ sentences when more than $k$ sentences are selected by the simulated knowledge selector***This random sampling does not break the experiment because our knowledge retriever and selector are simulated via random sampling, which is not the case for actual model-based retrievers and selectors.. We choose $k=3$ for both WoW and HotpotQA because their average number of candidate knowledge sentences before random sampling are 7.1 and 9.5 respectively.
@@
Figure 9: Scatter plot of answer F1 versus the knowledge precision for GPT-4o-mini on Wizard of Wikipedia (left) and HotpotQA (right) by limiting the number of knowledge sentences up to k=3. The dots highlighted in orange indicate settings outperforming the “full knowledge” setting (without constraining the number of sentences for fair comparison with other figures), while those highlighted in cyan indicate settings underperforming the “no knowledge" setting. Each data point corresponds to a full experiment on the entire sampled dataset.

### v3

@@ 6.2 Recommendations to Practitioners
Only when the knowledge recall score is high can a knowledge selector, which increases knowledge precision, be potentially helpful [^2] In a length-constrained scenario, if we consider $k$ randomly sampled knowledge sentences as a baseline, rather than the “full knowledge” setting, then the knowledge selector is more likely to be helpful due to weaker baseline performance.. Moreover, if the recall is too low, the downstream generation performance may not see any benefits from a knowledge selector with middling performance — it may even be harmed; only when the knowledge selector’s precision is very high will the overall performance improve.
@@
Due to the high cost of larger-scale simulations in each meta-experiment that consists of hundreds of full experiments using API-based LLMs, we use only a subset of the WoW and HotpotQA for experiments. As a result, our data subset may contain some minor noise during knowledge sampling and LLM generation, and the contours in Figure 4 are not smooth. However, such noise is not likely to affect our conclusions. For a similar reason, we only chose three LLMs for our experiments. As a result, we may have missed out on more subtle phenomena that can only be seen from results on a larger number of generators.
@@ A.1 Implementation Details
A.1.1 Knowledge Sampling
A.1.2 Models
We do not perform any fine-tuning or hyperparameter tuning. We set the LLMs’ temperature to 0 and ensure that our zero-shot generation prompts (Appendix A.2) are shorter than their maximum input lengths. We use the API services of OpenAI gpt-4o-mini-2024-07-18333https://platform.openai.com/docs/models/gpt-4o-mini, Together AI [^4] https://docs.together.ai/docs/chat-models meta-llama/Meta-Llama-3.1-8B-Instruct-Turbo, and mistralai/Mistral-7B-Instruct-v0.1. We report results on the first 500 examples in the HotpotQA training set and the first 100 conversations (452 wizard utterances) from the “test seen” set of WoW. Our set of experiments cost about 50 USD from OpenAI and about 50 USD from Together.ai.
@@ A.6 Length-Constrained Knowledge Selection and Answer Generation
To study the effect of length-constrained knowledge inputs on answer generation performance, we limit the number of knowledge sentences to $k$ by randomly sub-sampling $k$ sentences when more than $k$ sentences are selected by the simulated knowledge selector [^5] This random sampling does not break the experiment because our knowledge retriever and selector are simulated via random sampling, which is not the case for actual model-based retrievers and selectors.. We choose $k=3$ for both WoW and HotpotQA because their average number of candidate knowledge sentences before random sampling are 7.1 and 9.5 respectively.
@@
Figure 9: Scatter plot of answer F1 versus the knowledge precision for GPT-4o-mini on Wizard of Wikipedia (left) and HotpotQA (right) by limiting the number of knowledge sentences up to k=3. The dots highlighted in orange indicate settings outperforming the “full knowledge” setting (without constraining the number of sentences for fair comparison with other figures), while those highlighted in cyan indicate settings underperforming the “no knowledge" setting. Each figure is a meta-experiment, and each data point corresponds to a full experiment on the entire sampled dataset.